There are 27 nickels, dimes, and quarters in the drawer with a value of $3.00. How many coins of each type are there if there are three times as many nickels as there are dimes? Write and solve a system of equations.
Answer by stanbon(75887)

There are 27 nickels, dimes, and quarters in the drawer with a value of $3.00. How many coins of each type are there if there are three times as many nickels as there are dimes? Write and solve a system of equations.
---
Equations:
n + d + q = 27
5n +10d +25q = 300
n - 3d + 0q = 0
-----------------------
Solve by any method you know to get:
n = 15
d = 5
q = 7
===========
Cheers,
Stan H.
===========